Show that the energy of a simple harmonic oscillator in the n = 2 state is 5ℏω/2 by substituting the wave functionψ2 = A(2αx2- 1)e-αx2/2 directly into the Schroedinger equation, as broken down in the following steps.
First, calculate dψ2/dx, using A, x, and α.
dψ2/dx = ..........................

Second, calculate d2ψ2/dx2, using A, x, and α.
d2ψ2/dx2 = .........................

Third, calculate α2x2ψ2 - d2ψ2/dx2, using A, x, and α.
α2x2ψ2 - d2ψ2/dx2 = .......................

Fourth, calculate (α2x2ψ2 - d2ψ2/dx2)/ψ2, using A, x, and α.
(α2x2ψ2 - d2ψ2/dx2)/ψ2 = ...........................

Finally, calculate E ≡ [(α2x2ψ2 - d2ψ2/dx2)/ψ2]ℏ2/(2m), using A, x, α, m, ω, and ℏ.
E ≡ .............................

please answer the question as requested  
Answer #1

The given wave function is
  \psi_2(x)=A(2a x^2 -1)e^{-\frac{ax^2}{2}}
where,
   a=\frac{m\omega}{\hbar}
And so, we have

Step 1 :

  \frac{\mathrm{d} \psi_2}{\mathrm{d} x}=\frac{\mathrm{d} }{\mathrm{d} x}\left [A(2a x^2 -1)e^{-\frac{ax^2}{2}} \right ]
  \Rightarrow \frac{\mathrm{d} \psi_2}{\mathrm{d} x}=A\left [4ax-ax(2a x^2 -1) \right ]e^{-\frac{ax^2}{2}}
  \Rightarrow \frac{\mathrm{d} \psi_2}{\mathrm{d} x}=A\left [5ax-2a^2 x^3 \right ]e^{-\frac{ax^2}{2}}
Step 2 :
   \Rightarrow \frac{\mathrm{d}^2 \psi_2}{\mathrm{d} x^2}=\frac{\mathrm{d} }{\mathrm{d} x}\left [A\left (5ax-2a^2 x^3 \right )e^{-\frac{ax^2}{2}} \right ]
   \Rightarrow \frac{\mathrm{d}^2 \psi_2}{\mathrm{d} x^2}=A\left [(5a-6a^2x^2)-ax\left (5ax-2a^2 x^3 \right ) \right ]e^{-\frac{ax^2}{2}}
   \Rightarrow \frac{\mathrm{d}^2 \psi_2}{\mathrm{d} x^2}=A\left [5a-6a^2x^2-5a^2x^2+2a^3 x^4 \right ]e^{-\frac{ax^2}{2}}
   \Rightarrow \frac{\mathrm{d}^2 \psi_2}{\mathrm{d} x^2}=A\left [5a-11a^2x^2+2a^3 x^4 \right ]e^{-\frac{ax^2}{2}}

Step 3 :
   a^2x^2 \psi_2 - \frac{\mathrm{d}^2 \psi_2}{\mathrm{d} x^2}=a^2x^2A(2ax^2-1)e^{-\frac{ax^2}{2}}-A\left [5a-11a^2x^2+2a^3 x^4 \right ]e^{-\frac{ax^2}{2}}
  \Rightarrow a^2x^2 \psi_2 - \frac{\mathrm{d}^2 \psi_2}{\mathrm{d} x^2}=A\left [a^2x^2(2ax^2-1)-5a+11a^2x^2-2a^3 x^4 \right ]e^{-\frac{ax^2}{2}}
   \Rightarrow a^2x^2 \psi_2 - \frac{\mathrm{d}^2 \psi_2}{\mathrm{d} x^2}=A\left [10a^2x^2-5a \right ]e^{-\frac{ax^2}{2}}
   \Rightarrow a^2x^2 \psi_2 - \frac{\mathrm{d}^2 \psi_2}{\mathrm{d} x^2}=5a \left (A\left [2ax^2-a \right ]e^{-\frac{ax^2}{2}} \right )
Step 4 :
   \frac{a^2x^2 \psi_2 - \frac{\mathrm{d}^2 \psi_2}{\mathrm{d} x^2}}{\psi_2}=5a \frac{\left (A\left [2ax^2-a \right ]e^{-\frac{ax^2}{2}} \right )}{A\left [2ax^2-a \right ]e^{-\frac{ax^2}{2}} }
  \Rightarrow \frac{a^2x^2 \psi_2 - \frac{\mathrm{d}^2 \psi_2}{\mathrm{d} x^2}}{\psi_2}=5a
Step 5 :
   E=\left ( \frac{a^2x^2 \psi_2 - \frac{\mathrm{d}^2 \psi_2}{\mathrm{d} x^2}}{\psi_2} \right )\frac{\hbar^2}{2m}=5a \frac{\hbar^2}{2m}
  \Rightarrow E= 5\frac{\hbar^2}{2m}a
Now we use the definition of the parameter a
  a=\frac{m\omega}{\hbar}
And so,
   \Rightarrow E=5 \frac{\hbar^2}{2m} \times \frac{m\omega}{\hbar}
   \Rightarrow E=\frac{5}{2} \hbar \omega